# Qingyu Shi1  Lu Qi2  Jianzong Wu1   Jinbin Bai3

### v2

@@ 3.4.1 RelationLoRA Fine-tuning
where $Q=W_{q}h$, $K_{i}=W_{k_{i}}c_{i}$, $V_{i}=W_{v_{i}}c_{i}$, $K_{t}=W_{k_{t}}c_{t}$, $V_{t}=W_{v_{t}}c_{t}$, and $\gamma$ is a hyperparameter for scaling features from $c_{i}$. For clarity, we omit the final linear layer $W_{out}$. The predicate relation information mainly depends on text prompts. Therefore, we inject LoRA layers in all of the $W_{q}$, $W_{k_{t}}$, $W_{v_{t}}$, and $W_{out}$ to encourage the model to pay more attention to the predicate relation in $c_{t}$. We freeze all other parameters during fine-tuning. The required additional storage space for LoRA weights is only 13MB. As shown in Fig.4, benefited from the disentangled identity information in $c_{i}$, the LoRA weights effectively capture the predicate relation in $x$ and $c_{t}$, accurately generating the corresponding images. Notably, after fine-tuning, our RelationLoRA can also be directly integrated into SDXL (Podell et al., 2023), effectively enabling the model to address the relation inversion task. The extensive results are presented in the appendix (Fig. 15).
@@ 3.4.2 Keypoint Matching Loss
Most relations have specific requirements for the pose, such as “hugging” requires the arm to crossed over the other object’s body, and “riding a bicycle” requires the feet of the object to be on the pedal plate. Therefore, it’s important to accurately manipulate the pose of objects in the right status. Intuitively, we propose to introduce keypoint features into the fine-tuning procedure to explicitly guide the pose of the objects for better relation generation.
Specifically, considering our task involves common objects instead of humans only, We employ X-Pose (Yang et al., 2024) as keypoint detector because it can detect any keypoints in complex real-world scenarios. We detect 17 keypoints for each object in $x$ and $c_{i}$, and denote the keypoint coordinates as $c^{x}_{kp},c^{c_{i}}_{kp}\in\mathcal{R}^{17\times 2}$ respectively. To encourage the model to generate accurate pose in $\mathcal{D}(\hat{z}_{0})$, where $\mathcal{D}$ is the VAE decoder. We use the U-Net’s output $\hat{\epsilon}$ to predict $z_{0}$ during fine-tuning:
@@
We use the VAE encoder $\mathcal{E}$ to obtain latent representation of image prompts $\mathcal{E}(c_{i})$. Then we calculate the MSE loss on the corresponding keypoint locations between $\mathcal{E}(c_{i})$ and $\hat{z}_{0}$:
$\mathcal{L}_{KML}$ $=\frac{1}{n_{kp}}\mathbb{E}_{\bm{z_{t}},c_{i}}\left\|\mathcal{E}(c_{i})[c^{c_{i}}_{kp}]-\hat{z_{0}}[c^{z_{0}}_{kp}]\right\|^{2}_{2}$ (6)
@@
where $n_{kp}$ denote the number of keypoint, $c^{c_{i}}_{kp}$ and $c^{z_{0}}_{kp}$ are keypoint’s coordinates in $c_{i}$ and $z_{0}$ respectively, $\lambda$ controls for the relative weight of KML. The part inside the dotted box in Fig. 2 illustrates the model fine-tuning with the KML. We find KML is effective in encouraging the model to manipulate the pose for relation generation.
@@ 3.4.3 Local Token Injection
When generating relations, overlapping objects often occur. It is challenging to distinguish multiple objects solely using CLIP’s image-level embedding of the full image, which can lead to confusion between customized objects during relation generation. To address this limitation, we introduce a local region representation of images. Considering the domain shift from the full image to local image regions, we employ CLIPSelf’s self-distillation methods (Wu et al., 2024b) to enhance the region-language alignment of local region representation. Specifically, in line with CLIPSelf, we extract dense features (Zhou et al., 2022) from a student CLIP model and derive local tokens through partitioning and pooling (He et al., 2017). We partition the full image into patches and feed them into a frozen CLIP model to obtain teacher tokens. Our approach minimizes the cosine similarity between local tokens and teacher tokens to improve the region-language alignment of local tokens. To formulate this process, we denote the hidden states of the second-to-last layer of CLIP as $h^{\prime}_{clip}$. We modify the last layer of CLIP $h_{clip}=\operatorname{ModifiedAttention}(h_{clip})$ to obtain the dense features:
$h_{tmp}=\operatorname{Proj}_{v}(\operatorname{norm}(h^{\prime}_{clip})),h_{tmp}=h^{\prime}_{clip}+\operatorname{Proj}_{out}(h_{tmp}),h_{tmp}=h_{tmp}+\operatorname{FFN}(h_{tmp})$ (8)
where $\operatorname{Proj}$ means the linear projection. We discard the class embedding $h_{clip}[0]$ and reshape the image embeddings $z^{\prime}[1:h\times w]$ into an $h\times w$ feature map. We partition the feature map into patches, from which we can get local tokens $tok_{l}$ by pooling. During inference, we simply concatenate the aligned local tokens with the image-level tokens and send them to the ID extractor which is a transformer-based architecture shown in Fig. 2 b). The query $q$ is obtained by concatenating class text embedding with the bounding box.
$q=q+\operatorname{GatedSelfAttn}([q,tok_{i},tok_{l}])$ (9)
Experimental results show that local token injection enhances identity preservation in evaluation metrics and visual effects as shown in Fig. 9.
@@ 4 Experiments
DreamBooth (SDXL) 27.8 18.2 74.2 62.6 24.3 16.2 67.8 57.2
Custom Diffusion 26.5 15.2 73.2 58.8 20.1 15.4 64.7 55.3
Cones-V2 24.4 13.5 72.1 57.2 21.3 15.2 64.3 54.2
ELITE 25.7 14.9 75.4 61.5
AnyDoor 24.5 14.7 77.4 62.2 21.6 14.9 69.7 59.8
BLIP-Diffusion 26.2 15.7 77.4 57.7
SSR-Encoder 25.5 15.9 80.4 59.4 24.2 14.6 72.1 56.2
MS-Diffusion 26.5 18.8 78.7 64.5 26.9 18.9 73.8 58.8
@@
Figure 6: Multi-object comparison with training-based and tuning-based methods: Compared to our base model, our method shows a clear advantage in relation-aware generation and avoiding object confusion in overlapping scenarios.
Table 2: DreamBench.
@@
DreamBooth (SDXL) 31.2 81.5 69.2
Custom Diffusion 28.4 77.2 66.8
Cones-V2 31.0 76.5 67.2
ELITE 29.8 77.4 62.5
AnyDoor 25.5 82.1 67.8
SSR-Encoder 30.8 82.1 61.2
MS-Diffusion 31.5 79.3 68.2
Ours 31.6 78.9 67.4
Table 3: M-CustomConcept101.
@@
DreamBooth (SDXL) 29.5 67.4 49.2
Custom Diffusion 28.1 66.2 48.8
Cones-V2 29.2 66.5 47.2
AnyDoor 20.2 72.1 51.2
SSR-Encoder 30.6 71.1 52.2
$\lambda$-eclipse 29.2 68.2 48.2
MS-Diffusion 28.0 70.2 51.2
Ours 29.4 71.4 52.3
Figure 7: Additional results of relation-aware generation.
@@
Figure 9: Ablation study of keypoint matching loss (KML) and local tokens: The images are generated using the same random seed. We use green and red boxes to highlight the main differences. The areas within the green boxes show the improvements with our proposed component.
Table 4: Ablation study on our proposed components
Method Single-object Multi-object
CLIP-T CLIP-R CLIP-I DINO CLIP-T CLIP-R CLIP-I DINO
w/o Relation-aware Data 28.4 16.9 78.2 64.4 26.2 16.8 75.3 59.7
w/o Local Token Injection 29.2 21.0 76.8 62.5 28.5 19.5 75.1 59.3
w/o Keypoint Matching Loss 28.3 17.9 77.8 62.1 27.4 18.2 75.2 61.2
Full Model 30.6 21.4 77.9 63.4 28.9 20.4 75.4 62.1
@@
Lambda Multi-object
CLIP-T CLIP-R CLIP-I DINO
$\lambda$=1e-2 27.5 19.3 72.6 59.6
$\lambda$=1e-3 28.9 20.4 75.4 62.1
$\lambda$=1e-4 26.8 18.1 73.9 60.4
Table 7: Ablation study on the number of local tokens (determined by partition size).
Num Local Tokens Multi-object
CLIP-T CLIP-R CLIP-I DINO
N=2$\times$2 28.3 19.4 75.3 60.1
N=4$\times$4 28.9 20.4 75.4 62.1
N=8$\times$8 27.9 18.1 74.1 59.4
Table 8: Ablation study on local token injection methods to ID extractor.
@@
Implementation Details. For relation-aware fine-tuning, we generate a tuning set with 4-6 samples per relation. We incorporate LoRA layers into all text cross-attention layers of the U-Net. We set the LoRA rank to $r=4$, the parallel cross-attention scaling factor to $\gamma=0.6$, and the keypoint matching loss weight to $\lambda=1e-3$. The model is fine-tuned for 500 steps, using 2 A100 GPUs, with a total batch size of 8, completing the process in 10 minutes. We use the Adam optimizer with a learning rate of $41e-$ 4 and no weight decay, resulting in a total of 3.1M trainable parameters. Note that our RelationLoRA only modifies the text cross-attention layers in SDXL, making it compatible with any SDXL-based models. Considering the strong identity preservation capabilities of MS-Diffusion, we implement our RelationBooth on MS-Diffusion. During fine-tuning and inference, we concatenate local tokens with image tokens. We put more self-distillation details of the Local Image Encoder in appendix (6.1). Additionally, as shown in appendix (Fig. 15), our RelationLoRA is compatible with SDXL for addressing the Relation Inversion task (Huang et al., 2023).
Evaluation. To evaluate relation-aware customized image generation, we propose RelationBench, consisting of 44 objects from DreamBench (Ruiz et al., 2023) and CustomConcept101 (Kumari et al., 2023), along with 25 predicate relations. The object categories include pets, plushies, toys, people, and cartoons. Using GPT-4, we generate 100 cases for single- and multi-object evaluations. Following previous works (Ruiz et al., 2023; Wang et al., 2024), we evaluate our method on three metrics: (1) Identity Preservation, which assesses the similarity between the generated images and the image prompts, using the CLIP image score and DINO score that calculate the cosine similarity between the class embeddings of images, referred to as CLIP-I and DINO, respectively; (2) Text Alignment, which evaluates how well the generated images align with the text prompts, using the CLIP image-text score, denoted as CLIP-T; and (3) Relation Alignment. We observed that nouns in the prompts can inflate CLIP-T scores, making them less accurate for evaluating relation generation. To address this, we extract relation predicate from the prompts using spaCy and calculate the CLIP image-text score using only the predicate, denoted as CLIP-R.
4.1 Main Results
We evaluate our method through both quantitative and qualitative results in single- and multi-object cases. In single-object cases, we primarily focus on the alignment of the object’s pose with the predicate relation in the text prompt, as well as the model’s ability to preserve identity. In multi-object cases, in addition to relation generation and identity preservation, we also need to ensure there is no confusion between the multiple objects.
@@ 4.1 Main Results
Quantitative results. We quantitatively compare our method with baseline models across three benchmarks: RelationBench, DreamBench, and multi-object cases in CustomConcept101. As shown in Tab. 1, our method demonstrates a clear advantage on RelationBench for single-object cases, particularly in the CLIP-T and CLIP-R metrics. However, it performs slightly lower in the CLIP-I metric, likely due to significant pose variations in the objects, which can negatively impact the CLIP-I score. For multi-object generation, our method surpasses other approaches across all evaluation metrics, which we attribute to effectively preventing object confusion, leading to a significant improvement in CLIP-I performance. Furthermore, as shown in Tab. 3, we evaluate single-object generation performance on DreamBench. Our method outperforms others in the CLIP-T metric while delivering competitive results in CLIP-I and DINO. To evaluate multi-object generation, we conduct experiments on multi-object cases from CustomConcept101, which we denote as M-CustomConcept101. As shown in Fig. 3, our model achieves the highest performance in the DINO score and ranks second in both the CLIP-T and CLIP-I metrics.
Qualitative Comparison. We conduct qualitative comparison experiments for relation-aware customized image generation on RelationBench. For single-object comparisons, as shown in Fig. 5, our generated images are more accurate than those produced by our base model, MS-Diffusion, and naturally reflect the predicate relations in the text prompts. Additional single-object comparison results are provided in the appendix (Fig. 12). In multi-object comparisons, Fig. 6 demonstrates that our method excels in generating multi-object relations, successfully avoiding object confusion while accurately representing the intended relations.
4.2 Ablation study and Analysis
RelationLoRA Fine-tuning. Instead of injecting LoRA layers and fine-tuning on our curated $(x,c_{i},c_{t})$, we fine-tune the text embedding $R^{*}$ using a set of relation-specific data $(x,c_{t})$, where blank images are used as image prompts $c_{i}$. The visual comparison is presented in Fig. 4, and the quantitative results in Tab. 4 indicate that this method underperforms in both CLIP-T and CLIP-R scores. This may be due to the entanglement of identity and relation information in $x$, as well as the gap between the blank and real images, making it difficult to accurately generate the relation.
Keypoint Matching Loss. For the keypoint matching loss (KML), we omit it during fine-tuning. As shown in Fig. 9 a), including the KML leads to more accurate and natural relation-aware images when using the same random seed. Quantitative results, presented in Tab. 4, show a decline across all evaluation metrics when the KML is removed. Additionally, we experimented with different values for $\lambda$ and found that the best performance occurs at $\lambda=1e-3$.
Local Token Injection. For local token injection, the visual comparison in Fig. 9 b) shows severe object confusion when local tokens are omitted. The quantitative results in Tab. 8 demonstrate that local tokens improve relation-aware customization. We conduct further ablation as shown in Tab. 8, Tab. 6, and Tab. 8. We experimented with varying numbers of local tokens injected into the ID extractor and found that 16 tokens yielded the best results across all evaluation metrics. Therefore, we use 16 local tokens in all experiments. Additionally, we swapped the Local Image Encoder and found that CLIP-ViT-bigG was the most compatible model. For the injection method, we note that simply concatenating local tokens with image tokens produced the best outcomes. Combining these methods ensures that our method maintains high quality relation-aware customized generation.
5 Conclusion
In this work, we introduce a challenging task: relation-aware customized image generation, which aims to generate objects that adhere to the relations specified in the text prompt while preserving the identities of user-provided images. To support this, we propose a data engine for generating high-quality fine-tuning data. Our method, RelationBooth, incorporating Keypoint Matching Loss and Local Token Injection, effectively captures relation information and generates natural relations between customized objects while mitigating object confusion. We also provide a fair comparison with other methods on a new benchmark, RelationBench. The extensive experiments on three benchmarks demonstrate the effectiveness of our method, highlighting its potential for interactive scenario generation, relation detection, etc.
6 Appendix
This appendix provides additional details on the implementation of the Local Image Encoder and baseline methods. In the second section, we offer a brief introduction to the RelationBench benchmark. We then present further experimental results to validate the effectiveness of our method across various scenarios. Following that, we provide an in-depth analysis of the proposed components’ contributions to overall performance. Lastly, we discuss the potential ethical, privacy, and fairness implications of this technology, highlighting its social responsibility in real-world applications.
6.1 More Implementation Details

### v3

@@ 3.4.1 Relation Learning
where $Q=W_{q}h$, $K_{i}=W_{k_{i}}c_{i}$, $V_{i}=W_{v_{i}}c_{i}$, $K_{t}=W_{k_{t}}c_{t}$, $V_{t}=W_{v_{t}}c_{t}$, and $\gamma$ is a hyperparameter for scaling control from $c_{i}$. For clarity, we omit the final linear layer $W_{out}$. The relation information mainly depends on text prompts. Therefore, we inject LoRA layers in all of the $W_{q}$, $W_{k_{t}}$, $W_{v_{t}}$, and $W_{out}$ to encourage the model to pay more attention to the relationship in $c_{t}$. We freeze all other parameters during fine-tuning. As shown in Fig. 4, benefiting from the decoupled identity information in $c_{i}$, our method effectively captures the relationship in $x_{k}$ and $c_{t}$, accurately generating the corresponding images. Notably, after fine-tuning, our method can also be directly integrated into SDXL [27], effectively enabling the model to address the relation-aware generation task. The extensive results are presented in the Supplementary Material (SM).
Methods Single-Object Cus Multi-Object Cus Non-Confusing Relation-Aware
Textual Inversion [7]
DreamBooth [31]
Custom Diffusion [15]
Mix-of-Show [8]
CLIF [18]
MultiBooth [45]
BLIP-Diffusion [16]
ELITE [36]
AnyDoor [5]
SSR-Encoder [43]
$\lambda$-eclipse [24]
MS-Diffusion [35]
ReVersion [11]
ADI [10]
DreamRelation
Table 1: Setting Comparison for Different Models. We include several representative methods here. Our proposed DreamRelation can meet the demand of the relation-aware customization task.
Figure 4: Comparing our method with ReVersion across different base models, our approach demonstrates superior performance in the relation-aware generation task.
@@ 3.4.2 Keypoint Matching Loss
Most relationships have specific requirements for the pose, such as ”hugging” requiring the arms to cross over the object’s body, and ”riding a bicycle” requiring the feet of the object to be on the pedal plate. Therefore, it’s important to manipulate the pose of objects in the right status accurately. Intuitively, we introduce explicit supervision in the diffusion latent space during the tuning process, named Keypoint Matching Loss (KML).
Specifically, considering our task only involves common objects instead of humans, we employ X-Pose [41] as a keypoint detector because it can detect any keypoints in complex real-world scenarios. We detect 17 keypoints for each object in $x_{k}$ and $c_{i}$, and denote the keypoint coordinates as $c^{x_{k}}_{kp},c^{c_{i}}_{kp}\in\mathcal{R}^{17\times 2}$ respectively. To encourage the model to generate accurate poses in $\mathcal{D}(\hat{z}_{0})$, where $\mathcal{D}$ is the VAE decoder. We use the U-Net’s output $\hat{\epsilon}$ to predict $z_{0}$ during fine-tuning:
@@
We use the VAE encoder $\mathcal{E}$ to obtain the latent representation of image prompts $\mathcal{E}(c_{i})$. Then we calculate the MSE loss on the corresponding keypoint locations between $\mathcal{E}(c_{i})$ and $\hat{z}_{0}$:
$\mathcal{L}_{KML}$ $=\frac{1}{n_{kp}}\mathbb{E}_{\bm{z_{t}},c_{i}}\left\|\mathcal{E}(c_{i})[c^{c_{i}}_{kp}]-\hat{z}_{0}[c^{x_{k}}_{kp}]\right\|^{2}_{2}$ (6)
@@
where $n_{kp}$ denote the number of keypoint, $c^{c_{i}}_{kp}$ and $c^{x_{k}}_{kp}$ are keypoint’s coordinates in $c_{i}$ and $x_{k}$ respectively. We scale the coordinates to fit the size of $\mathcal{E}(c_{i}),\,\hat{z}_{0}$ $\lambda$ controls for the relative weight of KML. The part inside the dotted box in Fig. 2 illustrates the model fine-tuning with the KML. We find KML is effective in encouraging the model to manipulate the pose for relation-aware generation.
Figure 5: Additional results demonstrate the effectiveness of relation-aware generation. DreamRelation adapts the same pair of objects to different relationships in a natural and accurate manner.
@@ 3.4.3 Local Token Injection
When generating relationships, local and detailed features are essential—for instance, the feature of arms to accurately construct relationships like “carrying”. However, the coarse image-level features extracted by the CLIP Image Encoder fail to provide such fine-grained details, leading to confusion between objects in Fig. 7. To address this limitation, we introduce dense features from the CLIP Image Encoder [44]. Specifically, we modify the last layer of the CLIP Image Encoder to obtain dense features: $h_{\text{dense}}=\text{ModifiedAttention}(h^{\prime}_{\text{clip}})$, where $h^{\prime}_{\text{clip}}$ represents the hidden states of the second-to-last layer.
$h_{\text{tmp}}$ $=\text{Proj}_{v}(\text{norm}(h^{\prime}_{\text{clip}})),$ (8)
$h_{\text{tmp}}$ $=h^{\prime}_{\text{clip}}+\text{Proj}_{out}(h_{\text{tmp}}),$ (9)
$h_{\text{dense}}$ $=h_{\text{tmp}}+\text{FFN}(h_{\text{tmp}})$ (10)
To further enhance the compatibility between dense features and image-level features, we employ self-distillation on the CLIP Image Encoder [39] to derive our Local Image Encoder. Specifically, we align local regions of the dense features with their corresponding image-level features using cosine similarity.
During inference, we extract local tokens $tok_{\text{local}}$ from the dense feature $h_{\text{dense}}$ through partitioning and pooling. These local tokens are then concatenated with the image-level tokens $tok_{\text{image}}$ and passed to the ID extractor. The ID extractor, which is a transformer-based architecture as illustrated in Fig. 2 (b), can be written as:
$q=q+\text{Attention}(\text{concat}[q,\,tok_{\text{image}},\,tok_{\text{local}}])$ (11)
Regarding the compatibility between CLIP dense features and the original image-level features, experimental results demonstrate that our Local Image Encoder outperforms other dense representations [3]. For a detailed discussion, we refer readers to the Supplementary Material. Additionally, the injection of local tokens enhances identity preservation, as evidenced by improvements in evaluation metrics and visual results, shown in Fig. 7 and Tab. 3.
@@ 4 Experiments
Figure 6: We compare multi-object performance with training-based and tuning-based methods. Additionally, we integrate ReVersion [11] into MS-Diffusion, where the learnable text token is injected into the text prompt during MS-Diffusion’s inference, resulting in ReVersion+MS. Our method shows a clear advantage in relation-aware generation and avoiding object confusion in overlapping scenarios.
@@
DreamBooth (SDXL) [31] 27.8 18.2 74.2 62.6 24.3 16.2 67.8 57.2
Custom Diffusion [15] 26.5 15.2 73.2 58.8 20.1 15.4 64.7 55.3
Cones-V2 [20] 24.4 13.5 72.1 57.2 21.3 15.2 64.3 54.2
ELITE [36] 25.7 14.9 75.4 61.5
AnyDoor [5] 24.5 14.7 77.4 62.2 21.6 14.9 69.7 59.8
BLIP-Diffusion [16] 26.2 15.7 77.4 57.7
SSR-Encoder [43] 25.5 15.9 80.4 59.4 24.2 14.6 72.1 56.2
MS-Diffusion [35] 26.5 18.8 78.7 64.5 26.9 18.9 73.8 58.8
Reversion+MS [11] 27.8 19.3 77.8 63.1 27.2 19.2 74.2 59.7
@@
Figure 7: Ablation study of keypoint matching loss (KML) and local tokens: The images are generated using the same random seed. We use green and red boxes to highlight the main differences.
Figure 8: DreamRelation for three objects. Compared to the baseline model, DreamRelation shows a clear advantage.
Method Multi-object
CLIP-T CLIP-R CLIP-I DINO
w/o Relation-aware Data 27.3 19.4 75.3 59.8
w/o Local Token Injection 28.5 19.5 75.1 59.9
w/o Keypoint Matching Loss 27.4 19.2 75.2 61.2
Full Model 28.9 20.4 75.4 62.1
Table 3: Ablation study on our proposed components
Implementation Details. We generate a tuning set for relation learning with 4-6 samples per relationship. We incorporate LoRA layers into all text cross-attention layers of the U-Net. We set the LoRA rank to $r=4$, the parallel cross-attention scaling factor to $\gamma=0.6$, and the keypoint matching loss weight to $\lambda=1e-3$. The model is fine-tuned for 500 steps, using 2 A100 GPUs, with a total batch size of 8, completing the process in 10 minutes. We use the Adam optimizer [12] with a learning rate of $41e-$ 4 and no weight decay, resulting in a total of 3.1M trainable parameters. Note that our learnable parameters are only in the text cross-attention layers, making it compatible with any SDXL-based models. Considering the strong identity preservation capabilities of MS-Diffusion, we implement our DreamRelation on MS-Diffusion. During fine-tuning and inference, we concatenate local tokens with image tokens. We put more self-distillation details of the Local Image Encoder in the Supplementary Material.
Evaluation. To evaluate relation-aware customized image generation, we propose RelationBench, consisting of 44 objects from DreamBench [31] and CustomConcept101 [15], along with 25 relationships. The object categories include pets, plushies, toys, people, and cartoons. Using GPT-4, we generate 100 cases for single-object and multi-object evaluations. Following previous works [31, 35], we evaluate our method on three metrics: (1) Identity Preservation, which assesses the alignment between the generated images and the image prompts, using the CLIP image score and DINO score that calculate the cosine similarity between the class embeddings of images, referred to as CLIP-I and DINO, respectively; (2) Text Alignment, which evaluates how well the generated images align with the text prompts, using the CLIP image-text score, denoted as CLIP-T; and (3) Relation Alignment. We observed that nouns in the prompts can inflate CLIP-T scores, making them less accurate for evaluating relation generation. To address this, we extract the relation word from the text prompt using spaCy and calculate the CLIP image-text score, denoted as CLIP-R.
4.1 Main Results
We evaluate our method through both quantitative and qualitative results in single- and multi-object cases. In single-object cases, we primarily focus on the alignment of the object’s pose with the relationship in the text prompt and the model’s ability to preserve identity. In multi-object cases, in addition to relation generation and identity preservation, we also need to ensure no confusion between the multiple objects.
Quantitative results. We compare our method with baseline models across three benchmarks: RelationBench, DreamBench, and multi-object cases in CustomConcept101. As shown in Tab. 2, our method demonstrates a clear advantage on RelationBench for single-object cases, particularly in the CLIP-T and CLIP-R metrics. However, its performance is slightly lower in the CLIP-I metric, likely due to significant pose variations in the objects, which may decrease the CLIP-I score. For multi-object generation, our method outperforms other approaches across all evaluation metrics, which we attribute to its ability to effectively avoid object confusion and significantly enhance CLIP-I performance. Furthermore, as shown in Tab. 4, we evaluate single-object generation performance on DreamBench. Our method outperforms others in the CLIP-T metric while delivering competitive results in CLIP-I and DINO. We conduct experiments on multi-object cases from CustomConcept101, which we denote as M-CustomConcept101. As shown in Fig. 5, our model achieves the highest performance in the DINO score and ranks second in both the CLIP-T and CLIP-I metrics.
Qualitative Comparison. We conduct qualitative comparison experiments on RelationBench to evaluate relation-aware customized image generation. As shown in Fig. 6, our method outperforms others in generating multi-object relationships, effectively avoiding object confusion while accurately capturing the intended relationships. Additional results can be found in the SM.
4.2 Ablation study and Analysis
Relation-Aware Data Engine. Instead of fine-tuning on our curated $(x_{k},c_{i},c_{t})$ pairs. We train a text embedding $R^{*}$ using a set of relation-specific data $(x_{k},c^{\prime}_{i},c_{t})$, where $c^{\prime}_{i}$ is cropped from $x_{k}$. The well-trained text embedding $R^{*}$ is injected into the text embedding during inference. The quantitative results in Tab. 3 show that our relation-aware data engine outperforms in both CLIP-T and CLIP-R scores.
Keypoint Matching Loss (KML). We omit KML during fine-tuning. As shown in Fig. 7 a), including the KML leads to more accurate and natural relation-aware customized images using the same random seed. Quantitative results, presented in Tab. 3, show a decline across all evaluation metrics when the KML is removed. Additionally, we experimented with different values for $\lambda$ and found that the best performance occurs at $\lambda=1e-3$.
Local Token Injection. For local token injection, the visual comparison in Fig. 7 b) shows severe object confusion when local tokens are omitted. The quantitative results in Tab. 3 demonstrate that local tokens improve relation-aware customization. We conducted an ablation study on the number of local tokens, as presented in Tab. 6, and determined the optimal number to be 16. Additional ablation studies are shown in the SM. Combining these methods ensures that DreamRelation achieves high-quality relation-aware customized generation.
@@ 4.2 Ablation study and Analysis
DreamBooth (SDXL) [31] 31.2 81.5 69.2
Custom Diffusion [15] 28.4 77.2 66.8
Cones-V2 [20] 31.0 76.5 67.2
ELITE [36] 29.8 77.4 62.5
AnyDoor [5] 25.5 82.1 67.8
SSR-Encoder [43] 30.8 82.1 61.2
MS-Diffusion [35] 31.5 79.3 68.2
Ours 31.6 81.7 68.4
Table 4: Evaluation Results on DreamBench.
@@
DreamBooth (SDXL) [31] 29.5 67.4 49.2
Custom Diffusion [15] 28.1 66.2 48.8
Cones-V2 [20] 29.2 66.5 47.2
AnyDoor [5] 20.2 72.1 51.2
SSR-Encoder [43] 30.6 71.1 52.2
$\lambda$-eclipse [24] 29.2 68.2 48.2
MS-Diffusion [35] 28.0 70.2 51.2
Ours 29.7 71.4 52.3
Table 5: Evaluation Results on M-CustomConcept101.
Num Local Tokens Multi-object
CLIP-T CLIP-R CLIP-I DINO
N=2$\times$2 28.3 19.4 75.3 60.1
N=4$\times$4 28.9 20.4 75.4 62.1
N=8$\times$8 27.9 18.1 74.1 59.4
Table 6: Ablation study on the number of local tokens.
Lambda Multi-object
CLIP-T CLIP-R CLIP-I DINO
$\lambda$=1e-2 27.5 19.3 72.6 59.6
$\lambda$=1e-3 28.9 20.4 75.4 62.1
$\lambda$=1e-4 26.8 18.1 73.9 60.4
5 Conclusion
In this work, we introduce a new challenging task: relation-aware customized image generation. This task aims to generate objects that adhere to the relationship specified in the text prompt and preserve the identities of the user-provided images. To support this, we propose a data engine for generating high-quality fine-tuning data. Our method, DreamRelation, incorporates Keypoint Matching Loss and Local Token Injection, effectively capturing relation information and generating natural relations between customized objects. We also present a fair comparison with other methods on a new benchmark, RelationBench. Extensive experiments demonstrate the effectiveness of our method, highlighting its potential for applications in interactive scenario generation, relation detection, and more. Our research can inspire the direction of relation-aware generation in the community.
Acknowledgement. This work is supported by the National Key Research and Development Program of China (No. 2023YFC3807600).
Supplementary Material
Overview. The supplementary includes these sections:
  • Sec. 6. Introduction video.
  • Sec. 7. More experimental results.
  • Sec. 8. Implementation details of the experiments.
  • Sec. 9. Failure cases and limitations of our model.
  • Sec. 10. Details of the RelationBench.
  • Sec. 11. Incorporate with CogVideoX-I2V.
  • Sec. 12. Social impacts.
6 Introduction Video
We provide a video introduction of our work. Please refer to “introduction_video.mp4” in the supplementary file.
7 Additional Experimental Results
@@
Keypoint Matching Loss. We use X-Pose [41] as our keypoint detector due to its open-vocabulary detection capabilities that are compatible with a wide range of objects instead of humans only. The keypoint matching loss (KML) facilitates relation generation by explicitly guiding the model’s pose manipulation, resulting in more accurate pose generation. As shown in Fig. 11, the cat’s arm crosses the dog’s body, accurately depicting the ”hug” relation. The visualization in Fig. 10 shows the cosine similarity between the latent representation of the image prompt and model prediction. Implementation with KML shows a better alignment of specific parts during training and inference.
Figure 10: KML enhances the alignment of specific parts of the object in image prompt and model prediction.
Figure 11: Additional ablation studies on KML and Local Tokens.
Local Tokens Injection. To understand why local features enhance relation-aware generation, we adopt Principal Component Analysis (PCA) to project the dense feature more compactly. As shown in Fig. 12, dense features provide more fine-grained information than CLIP image tokens and provide detailed information about each object to construct the interaction in relation. Aiding in distinguishing between different objects during the generation process and helping to avoid object confusion, especially in cases of heavy overlap. Moreover, it can facilitate object appearance alignment. For the injection method, we note that simply concatenating local tokens with image-level tokens produced the best performance as shown in Fig. 9.
@@
Table 9: Ablation study on local token injection methods.
@@
Relation-Aware Customization Data Engine. Our relation-aware data are generated by DALLE-3 [2], which can effectively preserve the identity of some categories through its multi-turn dialogue capability. Specifically, we preserve the object’s identity by adding ”the photo of the same” in the text prompt. Although our relation-aware data only contained animal categories, it can effectively learn the relation information in images and generalize to a wide range of objects as shown in Fig. 9.
Figure 14: Our fine-tuning dataset as an example.
Relation-Aware Customized Image Generation. We show more qualitative results in Fig. 16, Fig. 21, and Fig. 9. We compare our methods with both training-based and tuning-based methods. Our method shows a clear advantage in pose manipulation to fit the relation. We conduct additional results in Fig. 15 to show DreamRelation’s differences with our base model-MS-Diffusion.
Relation Inversion Task. As shown in Fig. 17, our DreamRelation enhances SDXL’s ability to generate images that accurately adhere to specific relations. Compared to ReVersion [11], our method generates more precise relations, without object confusion or missing, demonstrating the strong performance of DreamRelation in the Relation Generation task.
Figure 17: Our DreamRelation is compatible with SDXL to address Relation Inversion task.
@@
Evaluation on DreamBench and CustomConcept101. As shown in Tab. 4 and Tab. 5, we evaluate single-object generation performance on DreamBench. Our method outperforms others in the CLIP-T metric while delivering competitive results in CLIP-I and DINO. To evaluate multi-object generation, we conduct experiments on multi-object cases from CustomConcept101, which we denote as M-CustomConcept101. As shown in Fig. 5, our model achieves the highest performance in the DINO score and ranks second in both the CLIP-T and CLIP-I metrics.
8 Additional Implementation Details